OpenStudy (anonymous):

Confirm my answer? I was asked to find dy/dx of x^2+y^2+3xy=9. I got -2x/(2y+3). Is this correct?

OpenStudy (amistre64):

x'^2+y'^2+3(xy)' = 9' 2x x'+ 2y y'+3(x'y+xy') = 0

OpenStudy (amistre64):

since dx/dx = x' = 1 2x + 2y y'+3(y+xy') = 0 2x + 2y y'+3y+3xy' = 0

OpenStudy (amistre64):

i like it

OpenStudy (amistre64):

(-2x-3y) up top

OpenStudy (amistre64):

2x + 2y y'+3y+3xy' = 0 2x + 3y + 2y y'+3xy' = 0 2x + 3y + (2y+3x)y' = 0

OpenStudy (anonymous):

\[F(x,y) = x ^{2}+y ^{2}+3xy-9=0\] \[F _{x}=2x+3y\] \[F _{y}=2y+3x\] \[dy/dx=-F _{x}/F _{y}=-(2x+3y)/(3x+2y)\]

OpenStudy (anonymous):

Thank you

OpenStudy (amistre64):

good luck ;)